OpenStudy (anonymous):

trains A and B are traveling in the same direction on the parallel tracks. train A is traveling 80 miles per hour and train B is traveling at 88 miles per hour. Train A passed a station at 12:20am if train B passes the same station at 12:35am At what time till train B catch up with train A?

OpenStudy (shubhamsrg):

when B reaches station,,A already travelled for 15 mins i.e 1/4 hrs distance = 1/4 * 80 =20 miles now for some time t , distance A travels = 80 t and for same time t , B travels = 88t B has to travel 20 miles more than A so as to meet it at some point so we have 80 t + 20 = 88t so calculate t and find out what time they actually meet.

OpenStudy (anonymous):

is it t=10?

OpenStudy (shubhamsrg):

t = 2.5 ..how you get 10 ?

OpenStudy (shubhamsrg):

2.5 hours *

OpenStudy (shubhamsrg):

and that 2.5 hrs is taken from 12:35 am ..was i clear ?

OpenStudy (shubhamsrg):

nevermind ^_^